OpenStudy (anonymous):

How do I solve these systems of equations x + y = 3 and x - y = 1 by graphing?

OpenStudy (anonymous):

well first you're going to want to get them in standard form.

OpenStudy (anonymous):

y=

OpenStudy (anonymous):

Oh... Y = mx + b? or just Y =?

OpenStudy (anonymous):

y=mx + b. Silly keyboard :D

OpenStudy (anonymous):

Once you get them into standard form, you'll be right next to graphing them (: I'll help with that too.

OpenStudy (anonymous):

IS y = x + 3 correct for the first one?

OpenStudy (anonymous):

For slope-intercept form (y = mx + b), you basically isolate the y.

OpenStudy (anonymous):

"IS y = x + 3 correct for the first one?" Close, you were suppose to subtract both sides by x.

OpenStudy (anonymous):

There I go again mixing up names again, thank you Micah. ><"

OpenStudy (anonymous):

You would subtract x from both sides, canceling the x on one side, but that negative should stick by aswell.

OpenStudy (anonymous):

Uhh.... y - x = 3?

OpenStudy (anonymous):

x + y = 3 -x -x

OpenStudy (anonymous):

So here, your x on the left should cancel and will end up on the right as a negative.

OpenStudy (anonymous):

solve for y on both equations and you will get y =3-x, intercept at (0,3) with slope of -1 and y= x-1, intercept at (0,-1), with slope positive 1

OpenStudy (anonymous):

Well, that did the work for him xD No good. ><"""

OpenStudy (anonymous):

^ Agreed, But don't worry I'm still working on understanding the process and not just slapping down the answers

OpenStudy (anonymous):

sometimes you learn faster knowing the answers first so you know what you need to do to get that answer

OpenStudy (anonymous):

Wait... Jayz said it's y = x - 1, I got -y = 1 - x

OpenStudy (anonymous):

all you have to do there is divide the equation by negative one, and you'll get that EXACT answer, you're on the right track and just a step away.

OpenStudy (anonymous):

Oh dividing it gives me -1 and makes the x positive, so i switch the order to y = x - 1 right?

OpenStudy (anonymous):

Yes! There you go (:

OpenStudy (anonymous):

Woo, thanks!

OpenStudy (anonymous):

Now how do I graph? :s

OpenStudy (anonymous):

Alright, you see that -1 there? That's going to be one of the intercepts. Do you know what intercepts are?

OpenStudy (anonymous):

Yes, so the -1 is the y intercept one right? so I put the point on -1 in the y line?

OpenStudy (anonymous):

Yeah, exactly!

OpenStudy (anonymous):

Once you have that there you're going to need to know it's slope, so you can add more 'points' and graph a line for that equation.

OpenStudy (anonymous):

Slope is 1 right? single x = 1?

OpenStudy (anonymous):

Yeah, the slope is positive and going up one. You're on the right track!

OpenStudy (anonymous):

Do you have an idea of how that would look?

OpenStudy (anonymous):

one up and one right or just one up?

OpenStudy (anonymous):

Ah, you're going to want to go one up and one right (:

OpenStudy (anonymous):

Oh yeah because 1/1 rise over run.

OpenStudy (anonymous):

and then I work on graphing y = 3 - x right?

OpenStudy (anonymous):

Yeah, exactly. I'll graph it for you, too. Just so you have an example. We still have the other one to do.

OpenStudy (anonymous):

Yeah, the same way the other one was done.

OpenStudy (anonymous):

Yeah mine looks like that

OpenStudy (anonymous):

Great, awesome. Good job!

OpenStudy (anonymous):

for y = 3 - x... 3 is the y intercept and 3 is the slope right?

OpenStudy (anonymous):

errr x is the slope*

OpenStudy (anonymous):

the y intercept is the 3, and your slope is the -x.

OpenStudy (anonymous):

so the points on this one are (0,3) and (-1, 2)?

OpenStudy (anonymous):

y=mx+b m is your slope b is your intercept In this case your m is -1x (: for future reference

OpenStudy (anonymous):

This one's slope isn't -1/2 though

OpenStudy (anonymous):

your -x is your slope and that's just like the last equation's slope x - but it's negative instead of positive.

OpenStudy (anonymous):

-1/1 right?

OpenStudy (anonymous):

Yeah!

OpenStudy (anonymous):

so the points are (0,3) (1,2)?

OpenStudy (anonymous):

exactly that. ^^

OpenStudy (anonymous):

You should have something like this.

OpenStudy (anonymous):

Exactly that, thanks a lot for helping me out with this question. Out of curiosity what do you use to make those graphs?

OpenStudy (anonymous):

MS paint with a background of a graph - nothing fancy at all. ^^ I wish I could tell you I used some kind of graphing program :P

OpenStudy (anonymous):

Great job, glad you were able to figure it out. (:

OpenStudy (anonymous):

Alright, seriously though thanks for the help, I was about to give up on this.

OpenStudy (anonymous):

O.o I'm glad you didn't Good luck with math!